OpenStudy (anonymous):

what is the value of y in the system of equations? 2x+6y=9 y=-1/2x+2

OpenStudy (anonymous):

the fraction threw me off & I can't find my calculator

OpenStudy (smurfy14):

x=3 y=1/2

OpenStudy (anonymous):

smurfy14 how did you get that?

OpenStudy (mattfeury):

kevork, this is your only warning to clean up your language. your account will be suspended.

OpenStudy (smurfy14):

substitute -1/2x+2 for the x in the first equation so it looks like this: 2x+6(-1/2x+2)=9 then distribute and solve for x. once you find x (which is 3) substitute in the 3 to the equation to make it look like this: 2(3)+6y=9 and solve from there and youll find y=1/2